# Dynamic Context Adaptation for Consistent Role-Playing Agents with Retrieval-Augmented Generations

### v3

@@ 6 Conclusion
In this work, we addresse critical limitations in building retrieval-augmented, RPAs with LLMs. By introducing a novel framework consisting of an Adaptive Context-aware Text Splitter (ACTS), Guided Selection (GS), and Attribute Extractor (AE), our approach enables robust and consistent simulation of character personas, even when confronted with queries that extend beyond explicit persona knowledge. Through the development of the CharacterRAG dataset, we provide a valuable resource for reproducible evaluation and benchmarking of RAG-based RPAs. Our experimental results demonstrate that the proposed method not only enhances the character’s knowledge representation, but also faithfully models nuanced traits such as personality. We are enthusiastic about the future prospects of RAG-driven role-playing agents, along with the creation of stronger character personas and improved RAG architectures.
@@
This study follows ethical guidelines to ensure the integrity and fairness of all experiments.
@@ Appendix C Related Work
Figure 5: AMADEUS framework. AMADEUS consists of three substage: (i) ACTS splits a persona document to make it suitable for RAG-based role-playing. (ii) To fully leverage the character’s knowledge, GS retrieves chunks from which it can infer the answer to the user query. (iii) AE uses the information derived from the GS results to extract character attributes.

### v4

@@ 6 Conclusion
In this work, we address critical limitations in building retrieval-augmented, RPAs with LLMs. By introducing a novel framework consisting of an Adaptive Context-aware Text Splitter (ACTS), Guided Selection (GS), and Attribute Extractor (AE), our approach enables robust and consistent simulation of character personas, even when confronted with queries that extend beyond explicit persona knowledge. Through the development of the CharacterRAG dataset, we provide a valuable resource that enables the construction of RAG-based RPAs, as well as their reproducible evaluation. Experimental results demonstrate that the proposed method not only enhances the character’s knowledge representation, but also faithfully models nuanced traits such as personality. We are enthusiastic about the future prospects of RAG-driven role-playing agents, along with the creation of stronger character persona dataset and improved RAG architectures.
@@
This work follows ethical guidelines to ensure the integrity and fairness of all experiments.
@@ Appendix C Related Work
Figure 5: AMADEUS framework. AMADEUS consists of three substages: (i) ACTS splits a persona document to make it suitable for RAG-based role-playing. (ii) To fully leverage the character’s knowledge, GS retrieves chunks from which it can infer the answer to the user query. (iii) AE uses the information derived from the GS results to extract character attributes.